OpenStudy (anonymous):

Given f(x)=[[3e^(1-(x^2))]*ln(x)] find the equation of the tangent line at x=1

OpenStudy (anonymous):

you know how to derivative?

sam (.sam.):

\[\frac{d}{dx}\left(\left(3 e^{1-x^2}\right) \ln (x)\right)=-\frac{3 e^{1-x^2} \left(2 x^2 \ln (x)-1\right)}{x}\] \[-\frac{3 e^{1-x^2} \left(2 x^2 \ln (x)-1\right)}{x}~,~x=1\] \[\frac{d}{dx}=3\] -------------------------------------------------------- y-y1=m(x-1)

sam (.sam.):

m=3

sam (.sam.):

use y-y1=m(x-1)

OpenStudy (anonymous):

Is the answer y=3x-3?

sam (.sam.):

yes